# ELIP: Enhanced Visual-Language Foundation Models for Image Retrieval

### v1

@@ 4.3 Training and Testing ELIP-B
Figure 3 illustrates the application of our architecture on BLIP-2. The only difference with CLIP/SigLIP is that BLIP-2 re-ranking is not via a dual encoder like CLIP/SigLIP, but the image and text encoders attend to each other. However, the purpose of our mapping network, and its training, are essentially unchanged.
Text-Guided Image-Text Matching Loss. At training time, we feed the text query ($T$) and the re-computed image features by the query text, i.e.,$\{v_{+},v_{-}\}$ ($v_{+}$ denotes the positive image and $v_{-}$ denotes the negative image), into the Q-Former, and then to an Image-Text Matching (ITM) Head to predict a score indicating whether the text and image match or not. The output of the ITM head is trained with binary cross entropy loss.
Re-Ranking at Inference Time. At inference time, for each text query, we first compute the similarity scores between the visual-language embedding, computed by the original BLIP-2 image and text encoders, to obtain an initial ranking of all images. We then select the top-$k$ candidates for further re-ranking, where, the visual features are re-computed by incorporating the prompted vectors from the mapping network. The final ranking is obtained via the sum of the initially computed similarity score and the score predicted by the ITM head based on the re-computed image features and text query.
@@ 5 Best Practice for Data Curation & Training
The recent visual-language foundation models are normally trained on huge numbers (billions) of paired image-caption samples, with considerable computing resources. In contrast, we explore a ‘student friendly’ best practice for data curation that enables improving large-scale visual-language models with limited resources. There are two major challenges: (i) training with a large batch size is challenging, due to limitation on GPU memory; (ii) it is difficult to train on billions of samples due to the prohibitive computation cost. Here, we discuss the strategy of global hard sample mining to make the training more effective with a small batch size in Section 5.1 and how to properly find and curate a large-scale image-text training dataset with maximum information gain in Section 5.2.
@@ 5.1 Global Hard Sample Mining
Training CLIP and BLIP-2 often requires a large batch size, as this increases the chances of having hard training samples, and leads to improved contrastive and discriminative capability of the model. Here, we adopt a strategy of global hard sample mining to group hard samples in batches to make training with small batch size more effective.
@@
Figure 4: Examples of generated training batches via global hard sample mining. For each row, the first sample is the one used to group other samples. Captions for row 1 (from left to right): a wooden table with no base; a wooden table with a couple of folding legs on it; a table that has a metal base with an olive wood top; small table outdoors sitting on top of the asphalt. Captions for row 2 (from left to right): a huge body of blue ice floats in a mountain stream; the big chunk of glacier is falling off of the cliff; there is a broken piece of glass that has been broken from the ground; a body of water surrounded by a forest near a mountain. It can be observed that the images and captions are very similar to each other, and significantly more close than images and captions in a random batch.
@@ 5.2 Selection and Curation of Large-Scale Dataset
In the literature, a number of large-scale image-text training datasets have been introduced, such as CC3M [69], DataComp [27], etc. A recent effort [78] utilises large-scale pre-trained image captioning models to generate synthetic captions for DataComp images, providing more information for training. Experiments [78] show that training CLIP on the generated DataCompDR12M dataset achieves better performance than on DataComp1B, although only 1% data samples are used. However, in our case, even using DataCompDR12M to train our models still takes prohibitive time to train ELIP-B on the 12M data with 2 A6000/A40 GPUs.
@@
Figure 5: Examples of the out-of-distribution benchmarks. Occluded COCO is on the left, and ImageNet-R is on the right. For both benchmarks, the positive images contain the object described by the text query while the negative images do not contain the object. We display positive images in the first row and negative images in the second row. For Occluded COCO, the target object in the image is always occluded, making it more difficult to be retrieved. For example, for the text query Bicycle in Occluded COCO, positive images have an occluded bicycle (highlighted in dashed box) while negative images do not have a bicycle in it; for the text query Goldfish in ImageNet-R, positive images have goldfish while negative images do not have goldfish.
Setting Architecture Training Dataset Hard Sample Mining Generate Multiple Prompts Recall@1 Recall@5 Recall@10 Average
$\mathbb{A}$ CLIP - 67.6 88.4 93.0 83.0
$\mathbb{B}$ ELIP-C CC3M [69] 68.8 88.9 93.8 83.8
$\mathbb{C}$ ELIP-C CC3M [69] 69.5 89.7 94.1 84.4
$\mathbb{D}$ ELIP-C DataCompDR [78] 71.3 90.6 94.4 85.4
$\mathbb{E}$ ELIP-C DataCompDR [78] 72.3 90.6 94.7 85.9
Table 1: Ablation study on Flickr. We use CLIP as the baseline for the ablation study as it is expensive to train BLIP-2.
@@
CLIP [65, 36] 2021 40.20 65.90 75.60 60.57 67.60 88.40 93.00 83.00
ELIP-C(Ours) - 45.60 71.10 80.40 65.70 72.30 90.60 94.70 85.87
SigLIP [93] 2023 47.80 72.40 81.00 67.07 74.70 92.30 95.60 87.53
ELIP-S(Ours) - 53.30 77.20 85.10 71.87 79.20 94.70 97.30 90.40
@@
Table 2: Comparison with recent state-of-the-art methods. Top: CLIP based models; Middle: SigLIP based models; Bottom: BLIP-2 based models. ELIP-C/ELIP-S significantly boosts the performance of CLIP/SigLIP, and ELIP-B outperforms the state-of-the-art BLIP-2 model. Results for models without * are zero-shot performance, whereas results for models with * are only zero-shot on Flickr as the BLIP-2 model has been fine-tuned on COCO and the * models are based on BLIP-2. However, our method brings an improvement over BLIP-2 on both benchmarks when trained on DataCompDR.
@@ 6.1 Standard Benchmarks
COCO is a large-scale object detection, segmentation, and captioning dataset. In terms of captioning, each image is annotated with 5 different captions. Previous works use the test split of 5,000 images and 25,010 captions for evaluation of text-to-image retrieval.
The Flickr30k Dataset contains images collected from Flickr, together with 5 reference sentences provided by human annotators. Its test set for text-to-image retrieval consists of 1,000 images and 5,000 captions.
Evaluation Measures. We use the standard measures for assessing retrieval performance: Recall@1, Recall@5 and Recall@10. Recall@k evaluates the proportion of relevant images that are successfully retrieved within the top-$k$ results for each text query, i.e., it indicates how often the ground-truth image appears in the top-$k$ ranked images returned by the retrieval algorithm, where the ground-truth image is the image that the query text caption is associated in the dataset.
@@ 6.2 Out-of-Distribution Benchmarks
To evaluate a model’s capability for text-to-image retrieval in out-of-distribution scenarios, we set up two new benchmarks for OOD text-based image retrieval, and they serve as complementary benchmarks to standard benchmarks such as Flickr and COCO. More specifically, Occluded COCO focuses on the retrieval of occluded objects, and ImageNet-R emphasizes retrieving objects from different unusual domains such as cartoons, sketches, etc.
Occluded COCO is generated using annotations from [44] with the method as described in [94] and focuses on the model’s capability of retrieving images with occluded target object against images that do not contain the target object.
ImageNet-R is generated using annotations from [33] and aims to examine the model’s performance in terms of ranking images of the target category in front of images of other categories on different domains, such as art, cartoons, deviantart, graffiti, embroidery, graphics, origami, paintings, patterns, plastic objects, plush objects, sculptures, sketches, tattoos, toys, and video game.
Figure 5 shows examples from the Occluded COCO and ImageNet-R benchmarks.
Evaluation Measures. Different from text-to-image retrieval using COCO / Flickr captions, on Occluded COCO and ImageNet-R there can be multiple positive images for each text query. Therefore, inspired by evaluation from the image retrieval community [62], we use mAP as the evaluation metric.
@@ 7 Experiment
Implementation Details. Due to the limitation of computing resources, we train the ELIP-C/ELIP-S models with a batch size of 40, and the ELIP-B model with a batch size of 12. We set the initial learning rate to be $31e-$ 3 to train ELIP-C/ELIP-S and $51e-$ 5 to train ELIP-B. We train models on DataCompDR by default, but also on smaller CC3M for an ablation study. Models are trained on 2 A6000 / A40 GPUs. In terms of re-ranking, for ELIP-C/ELIP-S, we re-rank the top-100 samples for COCO/Flickr, top-500 samples for Occluded COCO, and top-1000 samples for ImageNet-R; for ELIP-B, we re-rank the top-20 samples for COCO/Flickr, top-100 samples for Occluded COCO, and top-200 samples for ImageNet-R. The value of $k$ is chosen such that the recall at that $k$ is high in the original ranking, and also that the inference is fast. Please refer to the appendix for more details.
@@ 7.1 Results on COCO and Flickr Benchmarks
Ablation Study. In Table 1 we evaluate the benefits of different components of the ELIP framework for CLIP. Comparison between Settings $\mathbb{A}$ and $\mathbb{B}$ shows the effectiveness of the ELIP-C boost compared to the original CLIP. Comparing Settings $\mathbb{B}$ and $\mathbb{C}$ we can observe that it is essential to do hard sample mining when training the model at a small batch size. Comparing Settings $\mathbb{C}$ and $\mathbb{D}$ shows the benefit of training the model on a larger dataset with less noisy captions. Comparing Settings $\mathbb{D}$ and $\mathbb{E}$, it can be observed that it is benefitial to generate multiple visual prompts compared with generating 1 prompt. In the experiments here we are generating 10 prompts and we have conducted an ablation study in terms of the number of generated prompts in the appendix.
@@
Comparison with State-of-the-Art. Table 2 compares our models (ELIP-C, ELIP-S and ELIP-B) with previous models. It can be observed that when training on DataCompDR12M, our method achieves a zero-shot improvement for CLIP, SigLIP and BLIP-2 on COCO and Flickr benchmarks. Outperforming the most recent work [71], our ELIP-B therefore sets a new state-of-the-art for text-to-image retrieval on BLIP-2 backbone.
@@
Recall Top-$k$ Curves. Figure 6 shows the Recall Top-$k$ Curves of the original CLIP/SigLIP model and our ELIP-C/ELIP-S on the COCO benchmark. The Recall Top-$k$ curve is plotted via gathering the Recall value at different Top-$k$ values. We can note that there is a significant gap between the two models, indicating that ELIP-C/ELIP-S re-ranking is consistently improving text-to-image retrieval Recall over different K values.
@@
Figure 6: Recall Top-$k$ curve on COCO shows recall at different top-$k$ values for CLIP(left) and SigLIP(right) initial rankings and the re-rankings given by ELIP-C and ELIP-S, respectively.
@@
Figure 7: Qualitative comparison between the CLIP initial ranking and the ELIP-C re-ranking. COCO: Columns 1-2; Flickr: Columns 3-4. The positive image for each query is highlighted in the dashed box, and top-5 retrieved images are shown. For COCO, the text query is ‘A large wooden pole with a green street sign hanging from it’, CLIP ranks a pole not made of wood as top-1, while ELIP-C correctly re-ranks the large wooden pole back as top-1. For Flickr, the text query is ‘A man wearing bathing trunks is parasailing in the water’. CLIP ranks a wakeboarding person as top-1, while our ELIP-C model precisely re-ranks the parasailing person wearing bathing trunks at top-1.
@@
Qualitative Results. In Figure 7, we show a qualitative comparison between the initial ranking of the CLIP model and the re-ranking obtained after applying ELIP-C on COCO (left) and Flickr (right). The model significantly improves the ranking – in both cases the ground truth image (highlighted in the dashed box) is re-ranked from below the top-5 to be rank 1. Please refer to the appendix for more qualitative results.
@@
Figure 8: Visualisation of attention map shows the cross-attention map of [CLS] token on patch tokens for both CLIP and our ELIP-C. Left: image matches the text query; Right: image does not match the text query. For the situation where the image matches the text query, the difference in the attention maps shows our generated visual prompt vectors effectively improves the selection of image features relevant to the text query. For example, when the text query is ‘A young woman holding a giant tennis racket’ (Row 1), comparing the attention map of CLIP (Column 2) and ELIP-C (Column 3) we can observe ELIP-C enables more attention on the giant tennis racket and the young women; when the text query is ‘A baby in plaid shirt eating a frosted cake’ (Row 2), we note that ELIP-C (Column 3) gives significantly more attention on the cake, the baby and the shirt than CLIP (Column 2); when the text query is ‘A group of four women walking down a street in the rain’ (Row 3), it can be noted that ELIP-C (Column 3) largely increases the attention on the four women compared with CLIP (Column 2); when the text query is ‘A row of motorcycles parked in front of a building’ (Row 4), it can be noted that ELIP-C (Column 3) makes the attention to focus more on the row of motorcycles in comparison with CLIP (Column 2). The difference is not significant if the image does not match the query (Column 4-6).
@@
Visualisation of Attention Map. Figure 8 shows visualisations of the cross-attention map of the [CLS] token on patch tokens for both CLIP and our ELIP-C on COCO. It can be observed that our generated visual prompt vectors can effectively improve the selection of image features relevant to the text query if the image matches the text (left of Figure 8). As our ELIP-C fuses the text features at the very beginning of the image encoder, we hypothesize that this early fusion approach enables the model to generate image embeddings that are more closely aligned with the query text, thereby improving retrieval accuracy. The visualisations support this hypothesis.
@@
CLIP 50.9 76.0 63.5
ELIP-C (zero-shot) 52.4 77.0 64.7
ELIP-C (fine-tuned) 62.5 81.6 72.1
BLIP-2 62.7 82.3 72.5
ELIP-B (zero-shot) 63.4 83.0 73.2
ELIP-B (fine-tuned) 70.5 83.7 77.1
Table 3: Results on OOD datasets. Our method achieves a zero-shot improvement on both benchmarks and the peformance is further boosted when fine-tuned on COCO/ImageNet.

### v2

@@ 4.3 Training and Testing ELIP-B
Figure 3 illustrates the application of our architecture on BLIP-2. The only difference with that described for CLIP-type models is that BLIP-2 re-ranking does not use a dual encoder, rather, the image and text encoders attend to each other. However, the purpose of our mapping network, and its training, are essentially unchanged.
Text-Guided Image-Text Matching Loss. At training time, we feed the text query ($T$) and the re-computed image features with the query text as prompts, i.e.,$\{v_{+},v_{-}\}$ ($v_{+}$ denotes the positive image and $v_{-}$ denotes the negative image), into the Q-Former, and then to an Image-Text Matching (ITM) Head to predict a score indicating whether the text and image match or not. The output of the ITM head is trained with binary cross entropy loss.
Inference Time Re-Ranking. For each text query, we first compute the similarity scores between the visual-language embedding, computed by the original BLIP-2 image and text encoders, to obtain an initial ranking of all images. We then select the top-$k$ candidates for further re-ranking, where, the visual features are re-computed by incorporating the prompted vectors from the mapping network. The final ranking is obtained via the sum of the initially computed similarity score and the score predicted by the ITM head based on the re-computed image features and text query.
@@ 5 Best Practice for Data Curation & Training
The recent visual-language foundation models are often trained on massive numbers (billions) of paired image-caption samples, with considerable computing resources. Here, we explore a ‘student friendly’ best practice for data curation that enables improving large-scale visual-language models with limited resources. Specifically, there are two major challenges to be addressed: (i) training with a large batch size is challenging, due to limitation on GPU memory; (ii) training on billions of samples is prohibitively expensive on computation cost. In Section 5.1, we discuss the strategy of global hard sample mining to make the training more effective with a small batch size. In Section 5.2, we detail the procedure for selecting and curating an image-text training dataset with maximum information.
@@ 5.1 Global Hard Sample Mining
Training CLIP, SigLIP and BLIP-2 often requires a large batch size, as this increases the chances of having hard training samples, and leads to improved contrastive and discriminative capability of the model. Here, we adopt a strategy of global hard sample mining to group hard samples in batches to make training with small batch size more effective.
@@
Figure 4: Examples of generated training batches via global hard sample mining. For each row, the first sample is used to group other samples. Captions for row 1 (from left to right): ‘a wooden table with no base’; ‘a wooden table with a couple of folding legs on it’; ‘a table that has a metal base with an olive wood top’; ‘small table outdoors sitting on top of the asphalt’. Captions for row 2 (from left to right): ‘a huge body of blue ice floats in a mountain stream’; ‘the big chunk of glacier is falling off of the cliff’; ‘there is a broken piece of glass that has been broken from the ground’; ‘a body of water surrounded by a forest near a mountain’. It can be observed that the images and captions are very similar to each other, and significantly more close than images and captions in a random batch.
@@ 5.2 Selection and Curation of Large-Scale Dataset
In the literature, a number of large-scale image-text training datasets have been introduced, such as CC3M [72], DataComp [28], etc. A recent effort [82] utilises large-scale pre-trained image captioning models to generate synthetic captions for DataComp images, providing more information for training. Experiments [82] show that training CLIP on the generated DataCompDR12M dataset achieves better performance than on DataComp1B, although only 1% data samples are used. However, in our case, even using DataCompDR12M to train our models still takes prohibitive time (about 200 GPU days) to train ELIP-B on the 12M data with A6000/A40 GPUs.
@@
Figure 5: Examples of the out-of-distribution benchmarks. Occluded COCO is on the left, and ImageNet-R is on the right. For both benchmarks, the positive images contain the object described by the text query while the negative images do not contain the object. We display positive images in the first row and negative images in the second row. For Occluded COCO, the target object in the image is occluded, making it more difficult to be retrieved. For example, for the text query Bicycle in Occluded COCO, positive images have an occluded bicycle (highlighted in dashed box) while negative images do not have a bicycle in it; for the text query Goldfish in ImageNet-R, positive images have goldfish while negative images do not have goldfish.
Setting Architecture Training Dataset Hard Sample Mining Multiple Prompts COCO Avg. Flickr Avg.
R@1 R@5 R@10 R@1 R@5 R@10
$\mathbb{A}$ CLIP - 40.2 66.0 75.6 60.6 67.6 88.3 93.0 83.0
$\mathbb{B}$ ELIP-C CC3M [72] 40.7 66.2 76.1 61.0 68.8 88.9 93.8 83.8
$\mathbb{C}$ ELIP-C CC3M [72] 41.8 67.5 77.5 62.3 69.5 89.7 94.1 84.4
$\mathbb{D}$ ELIP-C DataCompDR [82] 44.2 70.0 79.5 64.6 71.3 90.6 94.4 85.4
$\mathbb{E}$ ELIP-C DataCompDR [82] 45.6 71.1 80.4 65.7 72.3 90.6 94.7 85.9
Table 1: Ablation study on ELIP-C for choice of training dataset, hard sample mining, and number of prompt vectors generated.
@@
CLIP [68, 37] 2021 40.16 65.95 75.62 60.58 67.56 88.34 93.00 82.97
ELIP-C(Ours) - 45.61 71.08 80.43 65.71 72.30 90.62 94.68 85.87
SigLIP [99] 2023 54.21 76.78 84.24 71.74 82.96 96.10 98.04 92.37
ELIP-S(Ours) - 61.03 82.62 88.70 77.45 87.62 98.16 99.16 94.98
SigLIP-2 [81] 2025 56.87 78.79 85.49 73.72 83.94 96.62 98.20 92.92
ELIP-S-2(Ours) - 62.91 83.86 89.70 78.82 87.74 97.96 98.94 94.88
@@
Table 2: Comparison with recent state-of-the-art methods. Top: CLIP-based models; Middle: SigLIP-based models; Bottom: BLIP-2-based models. ELIP-C/ELIP-S brings a significant zero-shot performance boost of CLIP/SigLIP architectures, and ELIP-B outperforms the state-of-the-art BLIP-2 model. Results for models without * are zero-shot, whereas results for models with * are only zero-shot on Flickr as the BLIP-2 model has been fine-tuned on COCO and the * models are based on BLIP-2. However, our method brings an improvement over BLIP-2 on both benchmarks when trained on DataCompDR.
@@ 6.1 Standard Benchmarks
COCO [50], is a large-scale dataset for studying object detection, segmentation, and captioning. In terms of captioning, each image is annotated with 5 different captions. Previous works use the test split of 5,000 images and 25,010 captions for the evaluation of text-to-image retrieval.
The Flickr30k Dataset [66] contains images collected from Flickr, together with 5 reference sentences provided by human annotators. The test set for text-to-image retrieval consists of 1,000 images and 5,000 captions.
Evaluation Metrics. We adopt the standard metrics for assessing retrieval performance, namely, Recall@1, Recall@5 and Recall@10. Recall@k denotes the proportion of relevant images that are successfully retrieved within the top-$k$ results for each text query.
@@ 6.2 Out-of-Distribution Benchmarks
To evaluate a model’s capability for text-to-image retrieval in out-of-distribution (OOD) scenarios, we set up two new benchmarks for text-based image retrieval. Specifically, Occluded COCO focuses on the retrieval of occluded objects, and ImageNet-R emphasizes retrieving objects from various unusual domains such as cartoons, sketches, etc. Figure 5 shows examples from the Occluded COCO and ImageNet-R benchmarks.
Occluded COCO is curated with annotations from [45], with the method as described in [100], where the occlusion relationship is utilised to collect images containing occluded objects. This dataset aims to evaluate the model’s performance on retrieving images with occluded target objects against images that do not contain the target object.
ImageNet-R is generated using annotations from [34] and aims to examine the model’s performance for retrieval across various domains, for example, art, cartoons, deviantart, graffiti, embroidery, graphics, origami, paintings, patterns, plastic objects, plush objects, sculptures, sketches, tattoos, toys, and video games.
Evaluation Metrics. Here, we use mAP as the evaluation metric. This is because there might be multiple positive images for each text query.
@@ 7 Experiment
Implementation Details. Due to computational resource constraints, we train the ELIP-C model with a batch size of 40, the ELIP-S model with a batch size of 10, and the ELIP-B model with a batch size of 12. The initial learning rate is set to $1\times 10^{-3}$ for ELIP-C, ELIP-S, and ELIP-S-2, and $1\times 10^{-5}$ for ELIP-B. All models are trained on the DataCompDR dataset by default, with additional experiments conducted on the smaller CC3M dataset for ablation studies. Training is performed on two A6000 or A40 GPUs. For re-ranking, we select the top-$k$ samples based on the dataset and model: for ELIP-C, $k$ is set to 100 for COCO and Flickr, 500 for Occluded COCO, and 1000 for ImageNet-R; for ELIP-S and ELIP-S-2, $k$ is set to 100 for COCO and Flickr, 500 for Occluded COCO, and 200 for ImageNet-R; for ELIP-B, $k$ is set to 20 for COCO and Flickr, 100 for Occluded COCO, and 200 for ImageNet-R. The value of $k$ is chosen to ensure high recall in the original ranking while maintaining fast inference. Compared to the original pre-training approach of CLIP, SigLIP and BLIP-2, our method significantly improves training efficiency in terms of reduced training time, GPU requirements, and batch size, with only a marginal increase in FLOPS introduced by the trainable MLP mapping network. Further details are provided in the appendix.
@@ 7.1 Results on COCO and Flickr Benchmarks
Ablation Study. In Table 1, we evaluate the contributions of different components of the ELIP framework for CLIP. A comparison between Settings $\mathbb{A}$ and $\mathbb{B}$ highlights the effectiveness of the ELIP-C boost over the original CLIP. The comparison between Settings $\mathbb{B}$ and $\mathbb{C}$ demonstrates the importance of hard sample mining when training with a small batch size. Settings $\mathbb{C}$ and $\mathbb{D}$ show the benefit of training on larger datasets with less noisy captions. Finally, the comparison between Settings $\mathbb{D}$ and $\mathbb{E}$ reveals that generating multiple visual prompts (e.g., 10 prompts in this study) is more beneficial than generating a single prompt. Further ablation studies on the number of generated prompts are detailed in the appendix.
@@
Comparison with State-of-the-Art. As shown in Table 2, we compare our models (ELIP-C, ELIP-S, ELIP-S-2, and ELIP-B) with prior state-of-the-art methods. When trained on DataCompDR12M, our method demonstrates zero-shot performance improvements for CLIP, SigLIP, SigLIP-2, and BLIP-2 on the COCO and Flickr benchmarks. Notably, ELIP-B outperforms the most recent work [74], establishing a new state-of-the-art for text-to-image retrieval on the BLIP-2 backbone. Furthermore, our ELIP-S, when applied to SigLIP and SigLIP-2, achieves performance comparable to BLIP-2.
@@
Recall Top-$k$ Curves. Figure 6 (right) presents the Recall@Top-$k$ curves for the original CLIP model and our ELIP-C on the COCO benchmark. The curves are generated by plotting the Recall values across various Top-$k$ thresholds. Notably, there is a significant performance gap between the two models, demonstrating that ELIP-C re-ranking consistently improves text-to-image retrieval performance across different $k$ values.
@@
Figure 6: Before/after comparisons. Left: Precision-Recall curves for Occluded COCO retrieval, comparing SigLIP-2 initial rankings to the re-rankings given by ELIP-S-2. Right: Recall Top-$k$ curves for COCO retrieval, comparing CLIP initial rankings to the re-rankings given by ELIP-C.
@@
Figure 7: Qualitative comparison between CLIP initial ranking and ELIP-C re-ranking. COCO: Columns 1–2; Flickr: Columns 3–4. The ground-truth image for each query is highlighted with a dashed box, with the top-5 retrieved images shown. For the COCO query “A large wooden pole with a green street sign hanging from it”, CLIP ranks a non-wooden pole as top-1, while ELIP-C correctly re-ranks the large wooden pole to top-1. For the Flickr query “A man wearing bathing trunks is parasailing in the water”, CLIP ranks a wakeboarding person as top-1, whereas ELIP-C accurately re-ranks the parasailing man wearing bathing trunks to top-1.
@@
Qualitative Results. Figure 7 provides a qualitative comparison between the initial rankings produced by the CLIP model and the re-ranked results obtained with ELIP-C on the COCO (left) and Flickr (right) benchmarks. In both cases, ELIP-C significantly improves the rankings by elevating the ground-truth image (highlighted with a dashed box) to rank 1. Additional qualitative results are provided in the appendix.
@@
Figure 8: Visualisation of attention maps comparing the cross-attention maps of the [CLS] token on patch tokens for CLIP and ELIP-C. Left: image matches the text query; Right: image does not match the text query. For matched queries, ELIP-C enhances attention on image features relevant to the text. For example: (Row 1) ELIP-C focuses more on the giant tennis racket and the young woman for the query “A young woman holding a giant tennis racket”; (Row 2) ELIP-C highlights the cake, baby, and shirt for the query “A baby in plaid shirt eating a frosted cake”. Differences are minimal when the image does not match the query (Columns 4–6).
@@
Visualisation of Attention Map. Figure 8 visualizes the cross-attention maps of the [CLS] token on patch tokens for both CLIP and ELIP-C on COCO. When the image matches the text query (left of Figure 8), our generated visual prompt vectors effectively enhance the selection of image features relevant to the query. This improvement can be attributed to ELIP-C’s early fusion approach, which integrates text features at the beginning of the image encoder, enabling the model to produce image embeddings more closely aligned with the query text. The visualizations provide strong evidence supporting this hypothesis.
@@
CLIP 47.47 76.01 61.74
ELIP-C (zero-shot) 48.89 76.81 62.85
ELIP-C (fine-tuned) 59.88 81.44 70.66
SigLIP 61.74 92.11 76.93
ELIP-S (zero-shot) 64.58 92.42 78.50
ELIP-S (fine-tuned) 71.99 92.86 82.43
SigLIP-2 66.40 92.66 79.53
ELIP-S-2 (zero-shot) 67.42 92.74 80.08
ELIP-S-2 (fine-tuned) 76.10 94.00 85.05
BLIP-2 62.73 82.31 72.52
ELIP-B (zero-shot) 63.40 82.99 73.20
ELIP-B (fine-tuned) 70.49 83.68 77.09
Table 3: mAP Results on OOD datasets. ELIP prompting achieves notable zero-shot improvements for CLIP, SigLIP series, and BLIP-2. These gains are further amplified through fine-tuning on relevant datasets. For example, to adapt to the Occluded COCO, the ELIP model is fine-tuned on COCO. Similarly, fine-tuning on ImageNet adapts ELIP to ImageNet-R. These results demonstrate ELIP’s capability for efficiently adapting the models to new datasets.